OpenStudy (anonymous):

Find the sum of the following infinite geometric series, if it exists. 2 + 1.5 + 1.125 + 0.8437 +…

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

s = a / 1-r = 2 /1 - .75 = 2 / 0.25 = 8

OpenStudy (anonymous):

ok? the common ratio r = 1.5 /2 = .75 and a = first term